OpenStudy (anonymous):

2. What is the value of cos 45°? (Points : 1) Option A: √2/2 Option B: √3/2 Option C: √2 Option D: √3

OpenStudy (anonymous):

\[\cos(45^\circ) = \frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

they're just equal.

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

oh cool i was confused on this one you shoiuld see my math work lol it fills up the whole page ^.^